[figure: 9.7 cm]

THE ANALOGY BETWEEN MAGNETOCALORIC BEHAVIOUR OF TYPE II. SUPERCONDUCTORS AND FERROMAGNETIC

MATERIAL

Z. Kalva Center for Theoretical Study – Charles University and Academy of Sciences, 11000 Praha 1, Czech Republic

The analogy between magnetocaloric behaviour of type II. superconductors and an anisotropic ferromagnetic material is discussed. The influence of anisotropy on the negative magnetocaloric effect near the phase transitions is illustrated by the numerical modelling.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : POSTER

Corresponding author : Z. Kalva

Address for correspondence : Center for Theoretical Study – Charles University and Academy of Sciences, 11000 Praha 1, Czech Republic

Email address : kalva@beba.cesnet.cz